a stream of nitrogen flowing at the rate of 2 kg/sec and the stream of hydrogen is flowing at the rate 0.5 kg/sec mixed adiabatically in a steady state flow process. if the gases are assumed ideal what is the rate of entropy change as a result of a process. in chemical engineering thermodynamics
1
Expert's answer
2020-03-23T10:14:41-0400

The entropy of adiabatic mixing n1 moles of an ideal gas 1 and n2 moles of an ideal gas 2 is given by the formula:

"\Delta S = R*(n1*ln((n1+n2)/n1)+"

"+n2*ln((n1+n2)\/n2))"

the moles of nitrogen N2 are: n1 = 2000 g/(28 g/mole) = 71.4286 mole

the moles of hydrogen H2 are: n2 = 500 g/(2 g/mole) = 250 mole

Putting n1 and n2 in the formula and multiplying by R = 8.31 J/(K*mole) gives that the rate of entropy increase per 1 second is:

"\Delta"S = 1414.883 J/(K*sec)